OpenStudy (anonymous):

What percent of 75 is 63? a119% b.47.25% c.94% d.84%

OpenStudy (anonymous):

Lets translate ... What can be translated as x and of as multiplication x * .........? What is next?

OpenStudy (anonymous):

x(75) ..... how do we translate "is"?

jhonyy9 (jhonyy9):

75 --- 100 63 ---- x -------------- x=63*100/75 = 63*4/3 = 21*4 =84

jhonyy9 (jhonyy9):

is 84%